# Information Re-Organization Improves Reasoning in Large Language Models

### v1

@@ 3 Methodology
3.1 Task Formulation
Given a sample $(\mathcal{C},\mathcal{Q},\mathcal{A})$, where $\mathcal{C}$ denotes gold context, $\mathcal{Q}$ denotes question, the task aims to obtain answer $\mathcal{A}$ to the question $\mathcal{Q}$, based on gold context $\mathcal{C}$ with a large language model. In particular, it requires multi-hop reasoning to get the answer $\mathcal{A}$ to question $\mathcal{Q}$.
3.2 Information Re-Organization
For a question $\mathcal{Q}$, if the derivation of its answer $\mathcal{A}$ relies on context $\mathcal{C}$, then a deeper understanding of $\mathcal{C}$ is crucial. Logical relationships, such as parallelism, causal connections, contrasts, etc., are essential elements of understanding and reasoning Braine (1978). However, the logical relationships in the original plain context are often implicit. Therefore, we uncover to obtain logical relationships in the context through information re-organization. MindMap Buzan et al. (2010) is a powerful structure for representing knowledge, concepts, and ideas, which can provide a hierarchical abstraction of plain text. Converting a text paragraph to this structure would provide an easier way to access the knowledge and ideas in the context. In addition, we choose this as the structure for re-organized information because it contains not only logical relationships but also multi-hop connections.
For example, for a multi-hop question: Where was the director of film The Green Fog born?, the corresponding context: The Green Fog is an experimental film directed by Guy Maddin that loosely revisits the plot of Alfred Hitchcock’s movie Vertigo through a collage of found footage … Guy Maddin (born February 28, 1956) is a Canadian screenwriter, director, author, cinematographer, and film editor of both features and short films, as well as an installation artist, from Winnipeg, Manitoba …, and the re-organized context is:
The Green Fog:
  Type: experimental film
  Director:
     Name: Guy Maddin
     Birth date: February 28, 1956
     Nationality: Canadian
     City: Winnipeg, Manitoba
     Profession: screenwriter, director
  plot: Alfred Hitchcock’s movie
[…]
Formally, MindMap elements are organized in hierarchical branches and reflect the logical relationships of elements. The same level can reflect parallel relationships and indented levels reflect causal relationships, as shown in Figure 2. In terms of content, MindMap is often around a central topic or idea. Based on the format and content criteria of the MindMap, we utilize the generative capabilities of large language models to construct it. According to the evaluation Chen et al. (2021), large language models are capable of generating information with a logical structure. Specifically, the large language model is asked to re-organize the context to a MindMap following instructions according to the context information and the question. Detailed information is shown in Appendix A.
3.3 Reasoning With Re-Organized Information
Following the re-organization process in Section 3.2, we get the re-organized context $\mathcal{G}$ from the context $\mathcal{C}$. Then the re-organized information $\mathcal{G}$ can be used as a context alone or combined with the original $\mathcal{C}$ to get the final answer of $\mathcal{Q}$. This enables LLMs to deeply understand the contextual content by clearly perceiving these logical relationships, facilitating the quality and reliability of reasoning.
Specifically, a large language model processes this context to infer the answer to question $\mathcal{Q}$ with a prompt. This step relies on the advanced natural language processing capabilities of large language models, encompassing understanding the context, integrating information, and logical reasoning based on this understanding. The output of a large language model is a response to question $\mathcal{Q}$, based on its comprehension of the context and analysis. The overall prompt template consists of three components: original context, e.g., documents or paragraphs, re-organized information, and a question. The prompt template is as follow:
Context:

### v2

@@ 3.3 Reasoning
4 Experimental
4.1 Tasks and Datasets
To verify the effectiveness of our information re-organization method, we conduct experiments across a range of contextually aware multi-hop reasoning tasks and datasets, including claim verification [12], question answering [13], and reading comprehension [14]. The detailed dataset information, including data splits and statistics, is available in Appendix D.
Claim Verification The task involves assessing a given claim against a set of evidence documents to determine whether they support or refute the claim [12]. We consider HOVER [24] and FEVEROUS [25], which comprise complex claims that necessitate multi-hop reasoning for verification. Besides, we also take into account the SCIFACT [26] dataset, notable for its inclusion of scientific claims.
Question Answering For this task, we consider the following datasets: 2WikiMultiHopQA [27], StrategyQA [28], MuSiQue [29], and HotpotQA [30]. To answer the questions in these datasets requires not only multi-hop reasoning but also cross-document analysis.
Reading Comprehension Machine reading comprehension task requires a model to process documents and select an answer from the provided candidates to a question about the content [14]. We primarily consider WIKIHOP [31] in the task, which necessitates multi-hop reasoning to derive the final answer. Additionally, HotpotQA [30] is also frequently considered as part of the QA domain.
4.2 Baselines
In our paper, we compare our method InfoRE to two reasoning baselines: 1) Standard. 2) CoT. The Standard approach is a method that directly reasons with the original textual context. The Chain of Thought (CoT) [6] method involves augmenting standard reasoning methods by adding a step-by-step thought process. In our paper, we conduct the CoT strategy by appending the sentence “Let’s think step by step.” at the end of the question.
The baseline methods and our InfoRE both adopt a zero-shot setting to counteract the potential randomness associated with demonstrations in a few-shot setting. We also design an answer-format instruction within the prompts for various tasks to standardize the structure of the final answer, thereby enhancing the precision of answer extraction. Moreover, all results reported in the paper use only the reorganized contextual information to reason. Comprehensive details about prompts and answer-format instruction are available in Appendix C. Following previous methods [17, 18, 25], we run the official evaluation scripts of each dataset to get the F1 to measure the results.
4.3 Implementation Details
In our paper, the large language models employed in the extraction and reasoning process include Llama2-70B [2], GPT-3.5 (text-davinci-003) [32] and GPT-4 [3]. We configure all models with top_p parameter as $1.0$ and temperature as $0.0$. In the policy model, we use the BERT-base version on all tasks and datasets. In RL training, we calculate the F1 score between the generated answer and the reference answer as the reward, with a rescaling coefficient of 10. We train the model for 1000 episodes. We conduct training for epoch 5, a batch size of 4, and a learning rate of 2e-6. The parameter of $\epsilon$ is set to 0.2. All experiments are conducted on an NVIDIA RTX A6000.
5 Results and Analysis
5.1 Main Results
Claim Verification
Table 1 presents a comprehensive performance comparison between our InfoRE and existing zero-shot techniques. For the HOVER dataset, we segment it into the 2-hop, 3-hop, and 4-hop levels following previous methods [17]. As depicted in Table 1, our InfoRE demonstrates significant improvements in the zero-shot claim verification task. The CoT [6] approach offers a lightly increase of 0.62% in the HOVER 4-hop using GPT-4, which indicates its marginal utility in more complex reasoning scenarios. Yet, our InfoRE achieves 3.02% improvement on the HOVER 4-hop using GPT-4 showing remarkable performance on contextual aware understanding and reasoning. This improvement is further increased to 73.62% in combination with CoT, suggesting that the methods complement each other effectively. In the case of Llama2-70B, the combined application of InfoRE and CoT yields a score of 53.20% on the 2-hop HOVER task, surpassing its CoT-only score of 50.02%. This pattern of improvement is consistent with GPT-3.5. GPT-4 shows superior performance across all methods and datasets, suggesting an inherent advanced reasoning ability. This performance is mirrored in the specialized benchmarks, where GPT-4 with InfoRE attains near-perfect accuracy on FEVEROUS (95.62%) and very high accuracy on SCIFACT (93.67%). These figures solidify the notion that the InfoRE indeed enhances the reasoning capabilities of LLMs.
LLMs Methods HOVER FEVEROUS SCIFACT
2-hop 3-hop 4-hop
LLAMA2-70B Standard 49.41 48.35 47.82 63.39 60.70
InfoRE 52.83 51.42 50.04 67.84 63.81
$\uparrow 3.42$ $\uparrow 3.07$ $\uparrow 2.22$ $\uparrow 4.45$ $\uparrow 3.11$
CoT 50.02 48.76 48.01 64.53 61.24
InfoRE + CoT 53.20 51.70 50.15 68.12 64.02
$\uparrow 3.18$ $\uparrow 2.94$ $\uparrow 2.14$ $\uparrow 3.59$ $\uparrow 2.78$
GPT-3.5 Standard 64.74 63.04 61.54 87.67 77.42
InfoRE 68.21 66.45 64.91 91.31 81.54
$\uparrow 3.47$ $\uparrow 3.41$ $\uparrow 3.37$ $\uparrow 3.64$ $\uparrow 4.12$
CoT 66.70 64.52 62.69 88.67 78.49
InfoRE + CoT 69.02 67.53 65.66 91.53 82.26
$\uparrow 2.32$ $\uparrow 3.01$ $\uparrow 2.97$ $\uparrow 2.86$ $\uparrow 3.77$
GPT-4 Standard 72.40 71.02 70.06 92.33 91.40
InfoRE 75.87 74.06 73.08 95.62 93.67
$\uparrow 3.47$ $\uparrow 3.04$ $\uparrow 3.02$ $\uparrow 3.29$ $\uparrow 2.27$
CoT 73.82 72.07 70.68 92.67 92.47
InfoRE + CoT 76.69 75.16 73.62 95.67 94.32
$\uparrow 2.87$ $\uparrow 3.09$ $\uparrow 2.94$ $\uparrow 3.00$ $\uparrow 1.85$
Question Answer and Reading Comprehension
Different from claim verification task, this task involves using multiple documents as context, presenting a challenge in cross-document reasoning. Intuitively, when the reasoning process involves multiple documents, information re-organization can effectively merge the information from different documents and uncover logical relationships that are not apparent in plain text. Performance in Table 2 verifies the intuitive. We can see that after applying information re-organization, all the results have a significant improvement. GPT-4 outperforms the other models, with the highest F1 score of 76.52%, 71.20%, and 83.22% when employing InfoRE on 2WikiMultiHopQA, StrategyQA, and HotpotQA, respectively. This is consistent with the performance improvements observed with GPT3.5 (text-davinci-003) and Llama2-70B when applying InfoRE. Different from conventional QA tasks, reading comprehension tasks require LLMs to not only deeply understand the context but also identify distractors among the candidates, increasing the reasoning challenge. Our InfoRE consistently shows improvements on this task. Specifically, in the WIKIHOP dataset, GPT-3.5 with InfoRE outperforms other methods with an F1 of 51.87%. This improvement further verifies the effectiveness of our method.
Table 2: Zero-shot results on Question Answering and Reading Comprehension tasks. 2WMHQA, SQA, and HQA are abbreviations for 2WikiMultiHopQA, StrategyQA, and HotpotQA, respectively.
LLMs Methods 2WMHQA MuSiQue SQA HQA WIKIHOP
LLAMA2 (70B) Standard 52.56 49.55 51.23 66.07 40.32
InfoRE 57.62 52.78 55.32 69.98 42.90
$\uparrow 5.06$ $\uparrow 3.23$ $\uparrow 4.09$ $\uparrow 3.91$ $\uparrow 2.58$
CoT 52.99 52.90 56.80 66.80 41.07
InfoRE + CoT 57.72 56.10 59.93 70.60 43.37
$\uparrow 4.73$ $\uparrow 3.20$ $\uparrow 3.13$ $\uparrow 3.80$ $\uparrow 2.30$
GPT-3.5 Standard 58.25 55.01 59.39 73.30 48.92
InfoRE 64.58 58.03 63.16 77.12 51.87
$\uparrow 6.33$ $\uparrow 3.02$ $\uparrow 3.77$ $\uparrow 3.82$ $\uparrow 2.95$
CoT 59.37 57.05 67.51 73.90 49.65
InfoRE + CoT 65.13 60.52 70.45 77.74 52.70
$\uparrow 5.76$ $\uparrow 3.47$ $\uparrow 2.94$ $\uparrow 3.84$ $\uparrow 3.05$
GPT-4 Standard 72.69 62.65 68.32 79.33 55.46
InfoRE 76.52 66.36 71.20 83.22 58.01
$\uparrow 3.83$ $\uparrow 3.71$ $\uparrow 2.88$ $\uparrow 3.89$ $\uparrow 2.55$
CoT 74.08 64.36 68.50 80.66 56.02
InfoRE + CoT 78.60 69.11 71.54 84.26 58.91
$\uparrow 4.52$ $\uparrow 4.75$ $\uparrow 3.04$ $\uparrow 3.60$ $\uparrow 2.89$
5.2 Analysis
Ablation studies
In our paper, the re-organization comprises two components: extraction and pruning. To investigate the impact of each component in detail, we conduct a series of ablation experiments using GPT-3.5 on the 2WikiMultiHopQA dataset. First, we directly remove extraction and pruning from our method, and the results are shown in the second and third rows of Table 3, respectively. It is worth mentioning that in the experiment where extraction is removed, we directly prune the sentences in the original context. Furthermore, we replace the reinforcement learning-based pruning method with a similarity-based pruning method to demonstrate its effectiveness. Specifically, the similarity-based pruning method uses Siamese-BERT, which takes the original question and each logical relationship as inputs separately, and then generates the corresponding representations. Then, we calculate the cosine similarity between these two representations. Finally, we removed the 30% of logical relationships with the lowest similarity, the result is shown in the last row of Table 3.
Table 3: F1 performance of ablation studies.
Methods 2WikiMultiHopQA
Full model 64.58
w/o extraction 61.64
w/o pruning 63.05
similarity-based pruning 63.32
The results in Table 3 show that removing the extraction and pruning operations leads to performance drops of 2.94% and 1.53%, respectively. This demonstrates the effectiveness of both components in our methods. The larger performance drop after removing extraction highlights the importance of extracting logical relationships for effective reasoning. After replacing the pruning model, the performance dropped by 1.26%, but the results were still better than without pruning. This not only demonstrates the necessity of pruning but also highlights the effectiveness of the reinforcement learning-based pruning method.
Quality of Re-Organized Information
To assess the quality of the re-organized information, we perform a quantitative evaluation of the re-organized information with GPT-4 (gpt-4-32k) on the 2WikiMultiHopQA dataset. Specifically, we select 100 samples from the dataset, GPT-4 (gpt-4-32k) is asked to rank re-organized information produced by GPT-3.5 (text-davinci-003) [1] and GPT-4, as well as original context following criteria: (1) Depth: The information present multiple relationships of a topic, offering insightful perspectives or in-depth understanding of the subject. (2) Clarity: Information is clear and precise, making it easy to understand without ambiguity. All of the information is ranked 1, 2, and 3 with 3, 2, and 1 scores, respectively. Finally, we get a weighted average score for each information to measure the overall quality.
Methods Depth
1st 2nd 3rd Avg R.
Original 0.22 0.36 0.42 1.80
GPT-3.5 0.32 0.36 0.32 2.00
GPT-4 0.46 0.28 0.26 2.20
Methods Clarity
1st 2nd 3rd Avg R.
Original 0.25 0.35 0.40 1.85
GPT-3.5 0.35 0.32 0.33 2.02
GPT-4 0.40 0.33 0.27 2.13
The results in Table 4 demonstrate that the re-organized information outperforms the original textual context in terms of depth and clarity, which justifies the motivation of our paper. In addition, the re-organized information from GPT-4 outperforms GPT-3.5, which proves in another way that GPT-4 is indeed more capable than GPT-3.5. The evaluation results further validate the effectiveness of our method. Furthermore, the 22.22% improvement in depth is more obvious than the 15.14% improvement in clarity, which indicates that the re-organization of the information has been particularly effective in enhancing the logical relationships of the information. The improvement in clarity suggests that the information is now presented in a more direct and streamlined manner, making it easier for LLMs to grasp the essential points without wading through unnecessary details.
Effect of Re-organized Information Quality
Methods FEVEROUS 2WikiMultiHopQA
GPT-3.5 (†)
    Standard 87.67 58.25
    InfoRE 91.31 64.58
    InfoRE * 92.50 66.61
GPT-4 (*)
    Standard 92.33 72.69
    InfoRE 95.62 76.52
    InfoRE 94.67 75.07
The cross-validation results, shown in Table 5, indicate that the use of either GPT-3.5 or GPT-4 for information re-organization leads to a marked improvement in the performance of LLMs on reasoning tasks. Additionally, it is observed that using GPT-4 for information re-organization while reasoning with the GPT-3.5 results in a further increase in reason results by 1.19% and 2.03% on FEVEROUS and 2WikiMultiHopQA datasets, respectively. This implies that enhancing the quality of re-organized information can improve the performance of LLMs, pointing to a promising direction for future research in refining information synthesis and re-organization. Conversely, when GPT-3.5 is employed for information re-organization in conjunction with reasoning using the GPT-4 model, there is a decrease in reason ability by 0.95% and 1.45% respectively. In addition, the magnitude of the decrease is lower than the increase, which implies that our re-organization strategy may have a greater impact on models with weaker reasoning abilities. This finding aligns with our understanding that models with inherently weaker reasoning abilities tend to rely more heavily on external strategy. For models with stronger inherent capabilities, our method still further improve its reasoning ability.
Error Analysis
To better comprehend where the errors in our InfoRE methodology come from and where they are fixed, we annotate 100 wrong predictions made by both InfoRE and Standard methods with GPT-3.5 on 2WikiMultiHopQA dataset. We categorize the errors into $4$ classes:
  • Contextual Misunderstanding (CM): This happens when the model fails to interpret or connect multiple pieces of information from different parts of the documents. Multi-hop reasoning requires synthesizing information from various segments, and recognizing logical relations, and any misunderstanding can lead to incorrect conclusion.
  • Factual Error (FE): The model may provide an answer that is factually incorrect or not supported by the given documents. This is often due to the model’s reliance on its training data, which may not always align with the specific facts in the context.
  • Mathematical Error (ME): The error occurs when math calculations are involved in deriving the final answer.
  • Unanswerable Question (UQ): It’s a specific type of error or limitation in dataset design, where the context does not contain enough information to provide a valid answer to the posed question.
Figure 4: Error Analysis of InfoRE on 2WikiMultiHopQA against Standard baseline method. The first four rectangles are error categories, while “Corrected" on the far right denotes the percentage of errors originally made by the baseline method that our method InfoRE has successfully corrected.
When engaging in reasoning with large language models, all four error categories are present. As shown in Figure 4, there are 6% unanswerable question errors in the dataset, and more than 90% errors occur in the reasoning in the baseline method. Among the four types of error, contextual misunderstanding is the primary source of errors in the baseline, highlighting the importance of an in-depth understanding of context in solving reasoning tasks with large language models. This finding is consistent with our motivation presented in the introduction section. Moreover, comparing the results of errors between our method and the baseline method, our method mainly corrects 14% of errors coming from the baseline method, most of the corrected errors are contextual misunderstanding errors. This further indicates that our InfoRE method assists Large Language Models (LLMs) in understanding context, signifying the necessity and effectiveness of conducting information re-organization before directly addressing the original question.
6 Conclusion
In this paper, we propose an information re-organization method to improve the reasoning ability of large language models (LLMs). Compared with previous approaches primarily focus on improving the quality of intermediate steps, our method emphasizes uncovering the logical relationships, multi-hop connections, and pruning the irrelevant information through information re-organization. This approach enables LLMs to explicitly perceive the logical relationships and multi-hop connections of concepts within the context, promoting a deeper integration and understanding of the context, which results in more robust reasoning outcomes. To verify the effectiveness of our method, we conduct experiments using various LLMs across a range of contextually aware multi-hop reasoning tasks. The experiment results demonstrate the potential of our method to improve the reasoning ability of LLMs. Additionally, our method has a positive impact on various tasks involving context understanding, such as academic research, legal analysis, and medical diagnostics. However, it is also important to be aware of potential negative impacts, such as the propagation of misinformation.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2024/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A Prompt for Extraction
The specific prompt of extraction is shown in Table  6. The purpose of the prompt is to obtain logical relationships and multi-hop connections from the context. In the prompt, the content within the brackets should be replaced with specific example content.
Table 6: Specific prompt for obtaining logical relationships. During execution, “[EVIDENCE]" needs to be replaced with the specific document, and “[CLAIM]" is replaced with the specific question.
Prompt Content for Logical Relationship Extraction
Given a claim and corresponding evidence, please summarize the evidence as a mind map according to the claim. The output must be in a strict JSON format: {“mind_map": “mind_map"}. CLAIM: [CLAIM] EVIDENCE: [EVIDENCE] .
Appendix B Input Format For Pruning
The extracted context $g$ includes logical relationships and corresponding attribute values. First, we iterate through all logical relationships and attribute values, and following the previous method [33], we concatenate them using [SEP] [CLS], combining them with the question to input into the pruning model. Then, we use the representation of the [CLS] token to represent the logical relationships for subsequent operations.
We use an example in Figure 2 to demonstrate the input format of the extracted context $g$ in the pruning model. The question is: Where did the producer of Julius Caesar study or work? The the extracted context $g$ is:
Julius Caesar:
  Production Company: Metro-Goldwyn-Mayer
  Director: Joseph L. Mankiewicz
  Producer:
     Name: John Houseman
     Education: Clifton College, England
     Occupation: Grain Trade, London
  Adaptation: Play by Shakespeare
We only traverse logical relationships of the first-level progressive type. Therefore, the logical relationships contained in extracted context $g$ of example include 1) Production Company: Metro-Goldwyn-Mayer, 2) Director: Joseph L. Mankiewicz, 3) Producer: Name: John Houseman, 4) Producer: Education: Clifton College,England, 5) Producer: Occupation: Grain Trade, London, 6) Adaptation: Play by Shakespeare. Then, we concatenate them using [SEP][CLS].
The input format after concatenation is: [CLS] Production Company: Metro-Goldwyn-Mayer [SEP] [CLS] Director: Joseph L. Mankiewicz [SEP] [CLS] [CLS] Producer: Name: John Houseman [SEP] [SEP] [CLS] Producer: Education: Clifton College [SEP] [CLS] Producer: Occupation: Grain Trade, London [SEP] [CLS] Adaptation: Play by Shakespeare [SEP] [CLS] Where did the producer of Julius Caesar study or work? [SEP].
Appendix C Prompt for Multi-hop Reason
During the reasoning stage using large language models, to accommodate more context-aware reasoning tasks while ensuring comparability of results, we designed a universal prompt template. The prompt template consists of three components: original context, e.g., documents or paragraphs, reorganized information, and a question. The prompt template is as follows:
Documents: